# Breaking the Factorization Barrier in Diffusion Language Models

### v1

@@ 1 Introduction
Diffusion language models (dLLMs) have recently emerged as a compelling paradigm for modeling natural language (Austin et al., 2021; Lou et al., 2024; Sahoo et al., 2024; Nie et al., 2025). Unlike traditional autoregressive language models that are bound to a fixed left-to-right generation order, dLLMs break this sequential constraint by offering the flexibility to make predictions in arbitrary orders and generate multiple tokens in parallel. By training a Transformer to predict distributions over sets of masked tokens simultaneously, dLLMs effectively enable global refinement of sequences, which offers a promising path toward bridging efficient parallel decoding with high-quality generation.
@@
The critical bottleneck in this formulation is the dimensionality of $\bm{\theta}$. Capturing even pairwise correlations would require parameters quadratic in the vocabulary size, which is computationally prohibitive. Consequently, ${p}_{\bm{\theta}}$ is typically restricted to fully factorized distributions, where $\bm{\theta}$ consists solely of independent logits for each position (Austin, 2018; Lou et al., 2024; Sahoo et al., 2024). This limitation forces a trade-off: highly dependent tokens must either be generated in separate steps or predicted simultaneously at the cost of incoherence.
@@ 5.1 Sampling Strategies
Latent Variable Sampling.   To circumvent the intractability of exact temperature scaling on the mixture distribution, we adopt an approximation based on the interpretation of PCs as deep latent variable models (Liu et al., 2023b; Peharz et al., 2016). Formally, the circuit defines a joint distribution $\hat{p}_{\bm{\theta},\bm{\omega}}(\bm{x})=\sum_{\bm{z}}\hat{p}_{\bm{\theta},\bm{\omega}}(\bm{x},\bm{z})$, where the discrete latent variables $\bm{z}$ represent the hierarchical routing decisions at sum nodes. The key insight is that conditioning on $\bm{z}$ resolves the intractable mixtures: once the latent path is fixed, the conditional distribution $\hat{p}_{\bm{\theta},\bm{\omega}}(\bm{x}|\bm{z})$ collapses into a single product term of leaf distributions. This allows us to apply temperature scaling exactly to the conditional component.
@@ 7 Related Work
Diffusion Language Models.   The adaptation of diffusion models to the discrete text domain began with D3PM (Austin et al., 2021), which formulated the forward process using stochastic transition matrices over categorical states. This foundation was extended by SEDD (Lou et al., 2024), which introduced a continuous-time framework to better model the discrete data distribution. A practical advancement came with Masked Diffusion Language Models (MDLM) (Sahoo et al., 2024), which derives a simplified training objective for masked diffusion models. Building on this masking paradigm, recent approaches like LLaDA (Nie et al., 2025) and Dream (Ye et al., 2025) have successfully scaled discrete diffusion to billions of parameters, demonstrating that non-autoregressive models can achieve competitive quality with autoregressive Transformers.
@@ Appendix D Justification of the Partition Function Computation Algorithm
Tractability. To establish the efficiency of this computation, we reuse Theorem 1 from Liu et al. (2023a).
@@ E.1 Latent-Space Modulation
The first approach leverages the interpretation of PCs as deep latent variable models (Peharz et al., 2016; Liu et al., 2023b). A PC can be viewed as a mixture model over a set of discrete latent variables $\mathbf{z}$, where each instantiation of $\mathbf{z}$ corresponds to a specific “routing path” through the sum nodes of the circuit. The joint distribution is given by $p_{\bm{\omega}}(\bm{x}_{0})=\sum_{\mathbf{z}}p_{\bm{\omega}}(\bm{x}_{0},\mathbf{z})$. Currently, the Transformer backbone $p_{\theta}$ only outputs potentials over the observed variables $\bm{x}_{0}$. We propose augmenting the Transformer to output a factorized distribution over both the data and the PC’s latent variables:

### v2

@@ 1 Introduction
Diffusion language models (dLLMs) have recently emerged as a compelling paradigm for modeling natural language (Austin et al., 2021; Lou et al., 2024; Nie et al., 2025). Unlike traditional autoregressive language models that are bound to a fixed left-to-right generation order, dLLMs break this sequential constraint by offering the flexibility to make predictions in arbitrary orders and generate multiple tokens in parallel. By training a Transformer to predict distributions over sets of masked tokens simultaneously, dLLMs effectively enable global refinement of sequences, which offers a promising path toward bridging efficient parallel decoding with high-quality generation.
@@
The critical bottleneck in this formulation is the dimensionality of $\bm{\theta}$. Capturing even pairwise correlations would require parameters quadratic in the vocabulary size, which is computationally prohibitive. Consequently, ${p}_{\bm{\theta}}$ is typically restricted to fully factorized distributions, where $\bm{\theta}$ consists solely of independent logits for each position (Austin, 2018; Lou et al., 2024; Shi et al., 2024; Sahoo et al., 2024). This limitation forces a trade-off: highly dependent tokens must either be generated in separate steps or predicted simultaneously at the cost of incoherence.
@@ 5.1 Sampling Strategies
Latent Variable Sampling.   To circumvent the intractability of exact temperature scaling on the mixture distribution, we adopt an approximation based on the interpretation of PCs as deep latent variable models (Liu et al., 2023; Peharz et al., 2016). Formally, the circuit defines a joint distribution $\hat{p}_{\bm{\theta},\bm{\omega}}(\bm{x})=\sum_{\bm{z}}\hat{p}_{\bm{\theta},\bm{\omega}}(\bm{x},\bm{z})$, where the discrete latent variables $\bm{z}$ represent the hierarchical routing decisions at sum nodes. The key insight is that conditioning on $\bm{z}$ resolves the intractable mixtures: once the latent path is fixed, the conditional distribution $\hat{p}_{\bm{\theta},\bm{\omega}}(\bm{x}|\bm{z})$ collapses into a single product term of leaf distributions. This allows us to apply temperature scaling exactly to the conditional component.
@@ 7 Related Work
Diffusion Language Models.   The adaptation of diffusion models to the discrete text domain began with D3PM (Austin et al., 2021), which formulated the forward process using stochastic transition matrices over categorical states. This foundation was extended by SEDD (Lou et al., 2024), which introduced a continuous-time framework to better model the discrete data distribution. A practical advancement came with Masked Diffusion Language Models (MDLM) (Sahoo et al., 2024) and MD4 (Shi et al., 2024), which derive a simplified training objective for masked diffusion models. Building on this masking paradigm, recent approaches like LLaDA (Nie et al., 2025) and Dream (Ye et al., 2025) have successfully scaled discrete diffusion to billions of parameters, demonstrating that non-autoregressive models can achieve competitive quality with autoregressive Transformers.
@@
Applications of Probabilistic Circuits and HMMs.   Prior research has demonstrated the utility of tractable probabilistic models in language tasks, including efforts to scale Hidden Markov Models for language modeling (Chiu and Rush, 2020) and utilizing tensor decompositions to improve multi-token prediction (Basharin et al., 2024). In controlled generation, PCs provide tractable guidance for both autoregressive language models (Zhang et al., 2024, 2023) and continuous diffusion models for tasks like image inpainting (Liu et al., 2024).
Most closely related to our hybrid parameterization are architectures like Conditional Sum-Product Networks (Shao et al., 2020) and Probabilistic Neural Circuits (Dos Martires, 2024), which leverage deep neural networks to directly output or condition the internal parameters of the tractable circuit. Instead of conditionally predicting all circuit parameters, which scales poorly to large vocabularies, CoDD efficiently captures dependencies by tractably multiplying a structural PC prior with a fully factorized neural output.
@@ Appendix D Justification of the Partition Function Computation Algorithm
Tractability. To establish the efficiency of this computation, we reuse Theorem 1 from Liu et al. (2024).
@@ E.1 Latent-Space Modulation
The first approach leverages the interpretation of PCs as deep latent variable models (Peharz et al., 2016; Liu et al., 2023). A PC can be viewed as a mixture model over a set of discrete latent variables $\mathbf{z}$, where each instantiation of $\mathbf{z}$ corresponds to a specific “routing path” through the sum nodes of the circuit. The joint distribution is given by $p_{\bm{\omega}}(\bm{x}_{0})=\sum_{\mathbf{z}}p_{\bm{\omega}}(\bm{x}_{0},\mathbf{z})$. Currently, the Transformer backbone $p_{\theta}$ only outputs potentials over the observed variables $\bm{x}_{0}$. We propose augmenting the Transformer to output a factorized distribution over both the data and the PC’s latent variables: